# {cutout}04.2cm55cm1 empty TextMonkey     : An OCR-Free Large Multimodal Model for Understanding Document

### 07 Mar 2024 v1

@@ Abstract
We present TextMonkey, a large multimodal model (LMM) tailored for text-centric tasks, including document question answering (DocVQA) and scene text analysis. Our approach introduces enhancement across several dimensions: by adopting Shifted Window Attention with zero-initialization, we achieve cross-window connectivity at higher input resolutions and stabilize early training; We hypothesize that images may contain redundant tokens, and by using similarity to filter out significant tokens, we can not only streamline the token length but also enhance the model’s performance. Moreover, by expanding our model’s capabilities to encompass text spotting and grounding, and incorporating positional information into responses, we enhance interpretability and minimize hallucinations. Additionally, TextMonkey can be finetuned to gain the ability to comprehend commands for clicking screenshots. Overall, our method notably boosts performance across various benchmark datasets, achieving increases of 5.2%, 6.9%, and 2.8% in Scene Text-Centric VQA, Document Oriented VQA, and KIE, respectively, especially with a score of 561 on OCRBench, surpassing prior open-sourced large multimodal models for document understanding. Code will be released at https://github.com/Yuliang-Liu/Monkey.
@@
Large Multi-modal Model, Document Analysis, Scene Text, Resolution
@@ 1 Introduction
Many early methods [1, 2] attempt to address the task using a two-stage approach: 1) detect and recognize the text using external systems. 2) document understanding based on the fusion of text results and images. However, the individual step of text reading in the processing pipeline may lead to the accumulation of errors. Moreover, relying on off-the-shelf OCR Models/APIs (OCR-Models) introduces additional engineering complexity, limits the connection between the text and its surrounding context, and can potentially increase computational costs. To alleviate the drawbacks of external systems before understanding, OCR-Free solutions [3, 4] have attracted increasing attention recently.
@@
The field of large multimodal models (LMMs) [5, 6] is advancing rapidly due to its powerful ability to handle diverse types of data. However, they still have limitations when it comes to addressing text-related tasks. As depicted in Fig. 1 (a), several methods, including LLaVAR [7], UniDoc [8], TGDoc [9], and mPLUG-DocOwl [10], heavily rely on a pre-trained CLIP [11] for visual encoding. Nevertheless, these encoders have input resolutions of 224 or 336, which are insufficient to meet the demands of documents containing numerous small texts [12]. Therefore, they can only recognize large text and struggle with small text in images. To address the limitations of tiny text, UReaer [13] and Monkey [14] take a cropping strategy to expand the input resolution, as shown in Fig. 1 (b). However, this crop strategy may inadvertently split related words, resulting in semantic incoherence. For example, the word "Backup" may be divided into "Back" and "up," making it impossible to restore its original meaning even after fusion has been performed. Besides, the spatial separation caused by this splitting also makes it challenging to handle text position-related tasks, such as text grounding. As shown in Fig. 1 (c), DocPedia [15] directly processes visual input in the frequency domain rather than the pixel space. Due to the characteristics of the frequency domain, it can quickly expand the resolution without losing information. However, due to the transformation of the feature space, it is difficult to leverage existing pretrained models, increasing the demand for training resources.
@@
We want to inherit the efficient image resolution scaling feature of Monkey [14] but address the missing cross-window context for the documents mentioned above. For this purpose, we introduce TextMonkey, as shown in Fig. 1 (d). TextMonkey utilizes a Split Module that divides high-resolution images into window patches using a sliding window method. Inspired by [16], we treat every self-attention layer in the CLIP as self-attention in non-overlapped windows. To introduce cross-window relationships while maintaining efficient computation, we use Shifted Window Attention with zero-initialization, which is similar to LoRA [17] to build cross-window connections. This approach allows us to maintain the training data distribution for the encoder and deal with high-resolution document images while reducing the computational cost of training from scratch. On the other hand, the utilization of the Split Module still poses a significant challenge as it leads to a notable increase in token length. We have observed that there are numerous repetitive image features that align with the language space, similar to certain repeated elements in the language itself. Thus, we propose a token resampler to compress these features while keeping as many of the most important features as possible. We employ important tokens as queries and the original features as key-value pairs, facilitating the reaggregation of features. On the basis of reducing the number of tokens, our module can also significantly improve the performance compared to random queries.
@@
Due to the self-explanatory nature of the text, in most cases, humans are able to locate the position of the answer itself. To alleviate the issue of hallucination in large language models further, we require the model not only to provide accurate answers but also to locate specific visual evidence supporting its response. We also introduce a variety of text-related tasks to deepen the connection between text information and visual information, such as text spotting and text grounding. Besides, incorporating positional cues into the answers can further enhance the model’s reliability and interoperability.
@@
  • Token compression. The enhancement in resolution results in some redundant tokens. By using similarity as a criterion, we are able to find significant tokens that serve as queries for the token resampler. This module not only reduces the token length but also improves the model’s performance. Additionally, it significantly improves the performance compared to the use of random queries.
  • Support text grounding. We expand our scope to include tasks beyond text-based question answering, encompassing reading text, text spotting, and text grounding. Additionally, we found that incorporating positional information into the answers not only improves the model’s interpretability but also minimizes instances of hallucination. TextMonkey can also be finetuned to understand the command of screen-shot clicking.
  • Our method inherently enhances many current evaluation datasets, resulting in average performance improvements with numerous baseline methods by 5.2%, 6.9%, and 2.8% for Scene Text-Centric VQA, Document Oriented VQA and KIE, respectively. Specifically, our model achieves a score of 561 on OCRBench, outperforming previous open-sourced large multimodal models (LMMs) designed for text understanding.
@@ 2 Related works
Models designed to comprehend images with text information can be broadly categorized into two types: OCR-Model-Driven and OCR-Free methodologies.
@@ 2.1 OCR-Model-Driven Methods
OCR-Model-Driven methods use OCR tools to acquire text and bounding box information. Subsequently, they rely on the models to integrate text, layout, and visual data. Meanwhile, diverse pre-training tasks are devised to enhance cross-modal alignment between visual and text inputs. StrucTexT [18] pays attention to the fine-grained semantic information and global layout information within the image in the design of pre-training tasks. Based on layout knowledge enhancement technology, ERNIE-Layout [19] innovatively proposes two self-supervised pre-training tasks: reading order prediction and fine-grained image-text matching. The LayoutLM [20, 21, 2] series continuously improves by integrating pre-trained text, layout, and visual features and introducing a unified model architecture and pre-training goals. This enhances the model’s performance in various document understanding tasks and simplifies overall design. UDOP [1] unifies vision, text, and layout through VTL Transformer and a unified generative pre-training task. Wukong-reader [22] proposes the Textline-Region Contrastive Learning and specially crafted pre-training tasks to extract fine-grained text line information. DocFormerv2 [23] designs an asymmetric pre-training method and a simplified visual branch for visual document understanding.
@@ 2.2 OCR-Free Methods
Although these methods do not require OCR tool limitations, they still need fine-tuning for specific tasks. In the fast-growing era of Multi-Modal Large Language Models (MLLMs), some models are explicitly trained on visual text understanding datasets and fine-tuned with instructions. LLaVAR [7], mPLUG-DocOwl [10] and UniDoc [8] create novel instruction-following datasets to enhance the tuning process and improve the comprehension of text-rich images. Additional efforts have been undertaken to capture more intricate textual details. UReader [13] designs a shape-adaptive cropping module that utilizes a frozen low-resolution visual encoder to process high-resolution images. DocPedia [15] processes visual input in the frequency domain rather than pixel space to process higher-resolution images with limited visual tokens. Recently, TGDoc [9] uses text-grounding to enhance document understanding.
Figure 2: An overview of the TextMonkey. It enables the enhancement of resolution with limited training resources while preserving cross-window information and reducing redundant tokens introduced by resolution enhancement.
@@ 3.1 Shifted Window Attention
To alleviate the issue mentioned above, we adopt Shifted Window Attention [16] to augment the CLIP model’s visual processing capabilities. Specifically, for an input image $I\in\mathbb{R}^{H\times W\times 3}$, our approach slices the image into non-overlapping windows. This slice is achieved using a sliding window $W\in\mathbb{R}^{H_{v}\times W_{v}}$, where $H_{v}$ and $W_{v}$ indicate the window’s size. Within each window, we independently apply a transformer block from the CLIP architecture, which initially does not account for cross-window relationships. To incorporate interactions between different windows and enhance the model’s contextual understanding of the image, we adopt the Shifted Window Attention (SWA) mechanism. The sliding window is cyclic-shifting toward the top-left direction, resulting in new windows. The self-attention computation by a masking mechanism, which limits self-attention computation to within new windows, as mentioned in  [16].
@@ 3.2 Token Resampler
As the resolution increases, the number of tokens also significantly increases. However, due to limitations in the input length of some language models and training time constraints, reducing the number of tokens becomes necessary. In common visual scenarios, the previous method [26] has demonstrated the feasibility of merging token approaches.
@@
However, how can we identify important tokens and eliminate redundant ones? We have observed that certain tokens are highly unique and lack closely similar counterparts, such as the fourth token in Fig. 3. This suggests that this token is distinct. We hypothesize that these tokens carry crucial and distinctive information, which is further validated in subsequent experiments. Therefore, we utilize similarity as a metric to identify significant tokens.
Hence, we propose a Token Resampler to compress redundant tokens, as shown in the left part of Fig. 2. Firstly, we utilize a token filter algorithm to select the most valuable tokens. The token filter algorithm is shown in Algor. 1.
@@ 3.3 Position-Related Task
However, this requires the model to have a strong spatial understanding and a better perception of the spatial positions of the text. Building upon the aforementioned model designs, we add additional training tasks to improve the model’s perception of text positions, such as text spotting and reading text. Specific tasks and prompts are shown in Tab. I. To ensure a robust association between text and location data, we strictly maintain their alignment, text information must precede any corresponding location details.
To standardize images of different ratios, we use a scale of (0, 1000) to represent positional information. Therefore, in an image with resolutions of ($H_{r}\times W_{r}$), the text coordinates (x, y) will be normalized to $[(x/H_{r}*1000)]$, and the same applies to y.
@@ 3.4 Dataset Construction
During our training process, we solely utilize open-source data and apply various task-specific augmentations to different datasets. By integrating various datasets and employing different instructions for different tasks, we enhance the model’s learning ability and training efficiency. For scene text scenario, we select COCOText [27], TextOCR [28], HierText [29], TextVQA [30], and MLT [31] for training. For document images, we select IIT-CDIP [32], DocVQA [33], ChartQA [34], InfoVQA [35], DeepForm [36], Kleister Charity (KLC) [37], and WikiTableQuestions (WTQ) [38]. The details of our training data are shown in Tab. II. We have a total of 409.1k pairs of dialogue data and 2.1M question-answering pairs in our dataset to train our model.
To further strengthen the model’s ability to handle structured text, we fine-tune on TextMonkey with structured data to enhance its structured capabilities, resulting in TextMonkey†. The fine-tuning data primarily consisted of 10% of the data from the previous stage, as well as a portion of structured data, including documents, tables, and charts. The structured data images are also sourced from publicly available datasets and are generated using their structure information.
@@ 4.1 Implementation Details
Model Configuration. In our experiments, we utilized the well-trained Vit-BigG and LLM from Qwen-VL [47], which is a pre-trained large multimodal model. We configured the height and width ($H_{v}$, $W_{v}$) of the image inputs to 448 to align with the encoder specifications of Qwen-VL. Our image resampler is equipped with 256 learnable queries, and the token resampler’s ratio (r) was set to 512 for images with a resolution of 896 and increased to 1024 for images with a resolution of 1344. To maximize training efficiency, our primary experimental focus was on outcomes at the 896 resolution setting.
Training. During the training phase, we utilized the AdamW [48] optimizer, setting the learning rate to 1e-5 for the initial stage and reducing it to 5e-6 for the subsequent stage, while adopting a cosine learning rate schedule. The parameters $\beta_{1}$ and $\beta_{2}$ were configured to 0.9 and 0.95, respectively. A warmup period comprising 150 steps was incorporated, and we processed the data in batches of 128. To mitigate the risk of overfitting, we applied a weight decay factor of 0.1. The comprehensive training procedure spanned across 12 A800 days to complete a single epoch.
@@
Table IV: Quantitative results on other document benchmarks.
Method Document Table Chart
DocVQA Deepform KLC WTQ ChartQA
Donut [3] 67.5 61.6 30.0 18.8 41.8
Pix2Struct [4] 72.1 - - - 56.0
UReader [13] 65.4 49.5 32.8 29.4 59.3
Qwen-VL [47] 65.1 3.1 13.9 21.6 65.7
Monkey [14] 66.5 40.5 33.9 25.3 65.1
TextMonkey 71.5 61.6 37.8 30.6 65.5
TextMonkey† 73.0 59.7 37.8 31.9 66.9

### 15 Mar 2024 v2

# TextMonkey: An OCR-Free Large Multimodal Model for Understanding Document

@@ Abstract
We present TextMonkey, a large multimodal model (LMM) tailored for text-centric tasks. Our approach introduces enhancement across several dimensions: By adopting Shifted Window Attention with zero-initialization, we achieve cross-window connectivity at higher input resolutions and stabilize early training; We hypothesize that images may contain redundant tokens, and by using similarity to filter out significant tokens, we can not only streamline the token length but also enhance the model’s performance. Moreover, by expanding our model’s capabilities to encompass text spotting and grounding, and incorporating positional information into responses, we enhance interpretability. It also learns to perform screenshot tasks through finetuning. Evaluation on 12 benchmarks shows notable improvements: 5.2% in Scene Text-Centric tasks (including STVQA, TextVQA, and OCRVQA), 6.9% in Document-Oriented tasks (such as DocVQA, InfoVQA, ChartVQA, DeepForm, Kleister Charity, and WikiTableQuestions), and 2.8% in Key Information Extraction tasks (comprising FUNSD, SROIE, and POIE). It outperforms in scene text spotting with a 10.9% increase and sets a new standard on OCRBench, a comprehensive benchmark consisting of 29 OCR-related assessments, with a score of 561, surpassing previous open-sourced large multimodal models for document understanding. Code will be released at https://github.com/Yuliang-Liu/Monkey.
@@
Large Multi-modal Model, Document Analysis, Scene Text, Resolution, OCRBench
@@ 1 Introduction
Many early methods [1, 2] attempt to address the task using a two-stage approach: 1) Detect and recognize the text using external systems; 2) Document understanding based on the fusion of text results and images. However, the individual step of text reading in the processing pipeline may lead to the accumulation of errors. Moreover, relying on off-the-shelf OCR Models/APIs (OCR-Models) introduces additional engineering complexity, limits the connection between the text and its surrounding context, and can potentially increase computational costs. To alleviate the drawbacks of external systems before understanding, OCR-Free solutions [3, 4] have attracted increasing attention recently.
@@
The field of large multimodal models (LMMs) [5, 6] is advancing rapidly due to its powerful ability to handle diverse types of data. However, they still have limitations when it comes to addressing text-related tasks. As depicted in Fig. 1 (a), several methods, including LLaVAR [7], UniDoc [8], TGDoc [9], and mPLUG-DocOwl [10] heavily rely on a pre-trained CLIP [11] for visual encoding. Nevertheless, these encoders have input resolutions of 224 or 336, which are insufficient to meet the demands of documents containing numerous small texts [12]. Therefore, they can only recognize large text and struggle with small text in images. To address the limitations of tiny text, UReaer [13] and Monkey [14] take a cropping strategy to expand the input resolution, as shown in Fig. 1 (b). However, this crop strategy may inadvertently split related words, resulting in semantic incoherence. For example, the word "Backup" may be divided into "Back" and "up," making it impossible to restore its original meaning even after fusion has been performed. Besides, the spatial separation caused by this splitting also makes it challenging to handle text position-related tasks, such as text grounding. As shown in Fig. 1 (c), DocPedia [15] directly processes visual input in the frequency domain rather than the pixel space. Due to the characteristics of the frequency domain, it can quickly expand the resolution without losing information. However, due to the transformation of the feature space, it is difficult to leverage existing pretrained models, increasing the demand for training resources.
@@
We want to inherit the efficient image resolution scaling feature of Monkey [14] but address the missing cross-window context for the documents mentioned above. For this purpose, we introduce TextMonkey, as shown in Fig. 1 (d). TextMonkey utilizes a Split Module that divides high-resolution images into window patches using a sliding window method. Inspired by [16], we treat every self-attention layer in the CLIP as self-attention in non-overlapped windows. To introduce cross-window relationships while maintaining efficient computation, we use Shifted Window Attention with zero-initialization to build cross-window connections. This approach allows us to maintain the training data distribution for the encoder and deal with high-resolution document images while reducing the computational cost of training from scratch. On the other hand, the utilization of the Split Module still poses a significant challenge as it leads to a notable increase in token length. We have observed that there are numerous repetitive image features that align with the language space, similar to certain repeated elements in the language itself. Thus, we propose a token resampler to compress these features while keeping as many of the most important features as possible. We employ important tokens as queries and the original features as key-value pairs, facilitating the reaggregation of features. On the basis of reducing the number of tokens, our module can also significantly improve the performance compared to random queries.
@@
On the other hand, due to the self-explanatory nature of the text, in most cases, humans are able to locate the position of the answer itself. To alleviate the issue of hallucination in large language models further, we require the model not only to provide accurate answers but also to locate specific visual evidence supporting its response. We also introduce a variety of text-related tasks to deepen the connection between text information and visual information, such as text spotting and text grounding. Besides, incorporating positional cues into the answers can further enhance the model’s reliability and interpretability.
@@
  • Token compression. We show enlarging resolution results in some redundant tokens. By using similarity as a criterion, we are able to find significant tokens that serve as queries for the token resampler. This module not only reduces the token length but also improves the model’s performance. Additionally, it significantly improves the performance compared to the use of random queries.
  • Support text grounding. We expand our scope to include tasks beyond text-based question answering, encompassing reading text, text spotting, and text grounding. Additionally, we found that incorporating positional information into the answers can improve the model’s interpretability. TextMonkey can also be finetuned to understand the command of screen-shot clicking.
  • We evaluated TextMonkey’s performance across 12 recognized benchmarks, observing significant improvements in several areas. Firstly, in scene text-centric tasks such as STVQA, TextVQA, and OCRVQA, TextMonkey achieved a 5.2% increase in performance. For document-oriented tasks, including DocVQA, InfoVQA, ChartVQA, DeepForm, Kleister Charity, and WikiTableQuestions, it showed a 6.9% improvement. In the domain of key information extraction tasks, like FUNSD, SROIE, and POIE, we noted a 2.8% uplift. Particularly notable was its performance in scene text spotting tasks (Total-Text, CTW1500, and ICDAR 2015) focused on transcription accuracy, where it improved by 10.9%. Additionally, TextMonkey set a new high score of 561 on OCRBench, a comprehensive benchmark encompassing 29 OCR-related evaluations, significantly surpassing the performance of previous open-source, large-scale multimodal models designed for document understanding. This achievement underscores TextMonkey’s effectiveness and advances in the field of document analysis and understanding.
@@ 2 Related works
Models designed to comprehend images with text information can be broadly categorized into two types: OCR-Model-Driven and OCR-Free methods.
@@ 2.1 OCR-Model-Driven Methods
OCR-Model-Driven methods use OCR tools to acquire text and bounding box information. Subsequently, they rely on the models to integrate text, layout, and visual data. Meanwhile, diverse pre-training tasks are devised to enhance cross-modal alignment between visual and text inputs. StrucTexT [17] pays attention to the fine-grained semantic information and global layout information within the image in the design of pre-training tasks. Based on layout knowledge enhancement technology, ERNIE-Layout [18] innovatively proposes two self-supervised pre-training tasks: reading order prediction and fine-grained image-text matching. The LayoutLM [19, 20, 2] series continuously improves by integrating pre-trained text, layout, and visual features and introducing a unified model architecture and pre-training goals. This enhances the model’s performance in various document understanding tasks and simplifies overall design. UDOP [1] unifies vision, text, and layout through VTL Transformer and a unified generative pre-training task. Wukong-reader [21] proposes the Textline-Region Contrastive Learning and specially crafted pre-training tasks to extract fine-grained text line information. DocFormerv2 [22] designs an asymmetric pre-training method and a simplified visual branch for visual document understanding. DocLLM [23] focuses exclusively on position information to incorporate the spatial layout structure, using a decomposed attention mechanism to build a cross-alignment between text and spatial modalities.
@@ 2.2 OCR-Free Methods
Although these methods do not require OCR tool limitations, they still need fine-tuning for specific tasks. In the fast-growing era of Multi-Modal Large Language Models (MLLMs), some models are explicitly trained on visual text understanding datasets and fine-tuned with instructions. LLaVAR [7], mPLUG-DocOwl [10] and UniDoc [8] create novel instruction-following datasets to enhance the tuning process and improve the comprehension of text-rich images. Additional efforts have been undertaken to capture more intricate textual details. UReader [13] designs a shape-adaptive cropping module that utilizes a frozen low-resolution visual encoder to process high-resolution images. DocPedia [15] processes visual input in the frequency domain rather than pixel space to process higher-resolution images with limited visual tokens. By training a visual vocabulary on a large amount of data, Vary [26] expands its resolution and achieves impressive results. Recently, TGDoc [9] uses text-grounding to enhance document understanding, suggesting that textual grounding can improve the model’s ability to interpret textual content, thereby enhancing its understanding of images rich in textual information.
Figure 2: An overview of the TextMonkey. It enables the enhancement of resolution with limited training resources while preserving cross-window information and reducing redundant tokens introduced by resolution enhancement. Besides, through various data and pretext prompts, TextMonkey has been equipped with the ability to handle multiple tasks.
@@ 3.1 Shifted Window Attention
To alleviate the issue mentioned above, we adopt Shifted Window Attention [16] to augment the CLIP model’s visual processing capabilities. Specifically, for an input image $I\in\mathbb{R}^{H\times W\times 3}$, our approach slices the image into non-overlapping windows. This slice is achieved using a sliding window $W\in\mathbb{R}^{H_{v}\times W_{v}}$, where $H_{v}$ and $W_{v}$ indicate the window’s size. Within each window, we independently apply a transformer block from the CLIP architecture, which initially does not account for cross-window relationships. To incorporate interactions between different windows and enhance the model’s contextual understanding of the image, we adopt the Shifted Window Attention (SWA) mechanism. As mentioned in  [16], the sliding window is cyclic-shifting toward the top-left direction, resulting in new windows. The self-attention computation by a masking mechanism, which limits self-attention computation to within new windows.
@@
3.2 Image Resampler
To reduce the redundancy in image features initially, we inherited the image resampler from Qwen-VL [28], which is using upon every window. The module employs a set of trainable parameters as query vectors and utilizes the image features from the visual encoder as keys and values for cross-attention operations. This process helps compress the visual feature sequence to a fixed length of 256. Furthermore, to preserve positional information crucial for fine-grained image comprehension, 2D absolute positional encodings are integrated into the query-key pairs of the cross-attention mechanism.
@@ 3.3 Token Resampler
As the resolution increases, the number of tokens also significantly increases, using the slide window mechanism. However, due to limitations in the input length of some language models and training time constraints, reducing the number of tokens becomes necessary. In common visual scenarios, the previous method [29] has demonstrated the feasibility of merging token approaches.
@@
However, how can we identify important tokens and eliminate redundant ones? We have observed that certain tokens are highly unique and lack closely similar counterparts, such as the fifth token in Fig. 3. This suggests that this token is distinct. We hypothesize that these tokens carry crucial and distinctive information, which is further validated in subsequent experiments. Therefore, we utilize similarity as a metric to identify significant tokens.
Hence, we propose a Token Resampler to compress redundant tokens, as shown in the left part of Fig. 2. As shown in Algor. 1, we utilize a token filter algorithm to select the most valuable tokens.
@@ 3.4 Position-Related Task
Moreover, we have undertaken modifications to existing question-answering datasets. Specifically, we have found the positions with the majority of answers in the images. These positional cues have been extracted and seamlessly integrated into the answers themselves. To preserve the original capability of direct dialogue, we have also retained the original question-answering task.
For better perception of the spatial positions of the text, it requires the model to have a strong spatial understanding. Building upon the aforementioned model designs, we add additional training tasks to improve the model’s perception of text positions, such as text spotting and reading text. Specific tasks and prompts are shown in Tab. I. To guarantee a strong connection between text and location data, we strictly maintain their alignment, ensuring that text information always comes before any associated location details.
To standardize images of different ratios, we use a scale of (0, 1000) to represent positional information. Therefore, in an image with resolutions of ($H_{r}\times W_{r}$), the text coordinates (x, y) will be normalized to $[(x/H_{r}*1000)]$, and the same applies to y. The restoration process involves the inverse operation.
@@ 3.5 Dataset Construction
During our training process, we solely utilize open-source data and apply various task-specific augmentations to different datasets. By integrating various datasets and employing different instructions for different tasks, we enhance the model’s learning ability and training efficiency. For scene text scenario, we select COCOText [30], TextOCR [31], HierText [32], TextVQA [33], and MLT [34] for training. For document images, we select IIT-CDIP [35], DocVQA [36], ChartQA [37], InfoVQA [38], DeepForm [39], Kleister Charity (KLC) [40], and WikiTableQuestions (WTQ) [41]. To accelerate the training speed, we have transformed single-image question answering into multi-turn image-based question answering, significantly improving the utilization of image features, following the successful approach introduced in LLaVA [5]. The details of our training data are shown in Tab. II. We have a total of 409.1k pairs of dialogue data and 2.1M question-answering pairs in our dataset to train our model.
To further strengthen the model’s ability to handle structured text, we fine-tune one epoch on TextMonkey with structured data to enhance its structured capabilities, resulting in TextMonkey†. The fine-tuning data primarily consisted of 5% of the data from the previous stage, as well as a portion of structured data, including documents, tables, and charts. The structured data images are also sourced from publicly available datasets and are generated using their structure information. Therefore, we have a total of 55.7k of data in structured data.
@@ 4.1 Implementation Details
Model Configuration. In our experiments, we utilized the well-trained Vit-BigG and LLM from Qwen-VL [28], which is a pre-trained large multimodal model. We configured the height and width ($H_{v}$, $W_{v}$) of the image inputs to 448 to align with the encoder specifications of Qwen-VL. Our image resampler is equipped with 256 learnable queries, and the token resampler’s ratio (r) was set to 512 for images with a resolution of 896 and increased to 1024 for images with a resolution of 1344. To maximize training efficiency, our primary experimental focus was on using TextMonkey and evaluating outcomes at the 896 resolution setting.
TextMonkey consists of a large language model with 7.7B parameters, an image resampler module with 90M parameters, a token resampler module with 13M, an encoder with 1.9B parameters, and Shifted Window Attention with 45M parameters. Overall, TextMonkey has a total of 9.7B parameters.
Training. During the training phase, we utilized the AdamW [50] optimizer, setting the learning rate to 1e-5 for the initial stage and reducing it to 5e-6 for the subsequent stage, while adopting a cosine learning rate schedule. The parameters $\beta_{1}$ and $\beta_{2}$ were configured to 0.9 and 0.95, respectively. A warmup period comprising 150 steps was incorporated, and we processed the data in batches of 128. To mitigate the risk of overfitting, we applied a weight decay factor of 0.1. The comprehensive training procedure spanned across 12 A800 days to complete one epoch.
@@
Table IV: Quantitative results on other document benchmarks. “DF” is an abbreviation for DeepForm.
Method Document Table Chart Scene
DocVQA DF KLC WTQ ChartQA TextVQA
Donut [3] 67.5 61.6 30.0 18.8 41.8 43.5
Pix2Struct [4] 72.1 - - - 56.0 -
UReader [13] 65.4 49.5 32.8 29.4 59.3 57.6
Qwen-VL [28] 65.1 3.1 13.9 21.6 65.7 63.8
Monkey [14] 66.5 40.5 33.9 25.3 65.1 67.6
TextMonkey 71.5 61.6 37.8 30.6 65.5 68.0
TextMonkey† 73.0 59.7 37.8 31.9 66.9 65.6